# A Kaczmarz-inspired approach to accelerate the optimization of neural network wavefunctions

### v1

@@ 3 Methods
Since $R$ is drawn from a continuous and high-dimensional space this equation can be seen as a highly overdetermined linear least-squares problem, with each configuration $R$ corresponding to a single row. We propose to solve this equation approximately by taking inspiration from the Kaczmarz method karczmarz1937angenaherte ; strohmer2009randomized . In particular, our strategy is to apply a single step of the randomized block Kaczmarz method needell2014paved to the previous parameter update $d\theta_{k-1}$ to obtain the new parameter update $d\theta_{k}$. This entails projecting $d\theta_{k-1}$ onto the solution space of $\bar{O}_{k}d\theta=\bar{\epsilon}_{k}$. Our algorithm thus differs from MinSR only in how we break the indeterminacy of the sampled subproblem: MinSR chooses the solution of minimal norm, whereas SPRING chooses the solution that is nearest to the previous parameter update.
A key assumption underlying this approach is that the imaginary time-step is small, meaning the parameter vector $\theta$ changes only slightly at each iteration. Thus, although the sampled subproblems $(\bar{O}_{k-1},\bar{\epsilon}_{k-1})$, $(\bar{O}_{k},\bar{\epsilon}_{k})$ can be completely different, the underlying overdetermined systems
$d\theta_{k-1}=\underset{d\theta}{\operatorname{argmin}}\;\mathbb{E}_{R\sim p_{k-1}}\left[\left|\bar{O}_{k-1}(R)d\theta-\bar{\epsilon}_{k-1}(R)\right|^{2}\right],$ (21)
$d\theta_{k}=\underset{d\theta}{\operatorname{argmin}}\;\mathbb{E}_{R\sim p_{k}}\left[\left|\bar{O}_{k}(R)d\theta-\bar{\epsilon}_{k}(R)\right|^{2}\right]$ (22)
are nearly identical. This enables us to understand the logic of SPRING as follows:
@@
    $d\theta_{k}=\underset{d\theta}{\operatorname{argmin}}\;\mathbb{E}_{R\sim p_{k}}\left[\left|\bar{O}_{k}(R)d\theta-\bar{\epsilon}_{k}(R)\right|^{2}\right].$ (23)
@@
    Apply the randomized block Kazcmarz method to these iterates to yield the SPRING update $d\theta_{k}$.
@@
Recall that the iterations of SPRING work by projecting the previous parameter update $d\theta_{k-1}$ onto the solution space of the current sampled subproblem $\bar{O}_{k}d\theta=\bar{\epsilon}_{k}$. With Tikhonov regularization, this amounts to the formula
@@
$d\theta_{k}=\underset{d\theta}{\operatorname{argmin}}\;\frac{1}{\lambda}\left\lVert\bar{O}_{k}d\theta-\bar{\epsilon}_{k}\right\rVert^{2}+\left\lVert d\theta-d\theta_{k-1}\right\rVert^{2}.$ (24)
@@
We find that directly using (24) results in unstable optimization (see Figure 7(b)). We do not yet have an explanation for this phenomenon. To stabilize the method, we decay the previous gradient by a small amount before projecting it. Formally, this stabilization comes in the form of a new regularization parameter $\mu$ and a modified update formula
@@
$d\theta_{k}=\underset{d\theta}{\operatorname{argmin}}\;\frac{1}{\lambda}\left\lVert\bar{O}_{k}d\theta-\bar{\epsilon}_{k}\right\rVert^{2}+\left\lVert d\theta-\mu d\theta_{k-1}\right\rVert^{2}.$ (25)
@@
An explicit formula for the parameter update $d\theta_{k}$ can now be derived. We first define $\phi_{k}=d\theta_{k}-\mu d\theta_{k-1}$, $\phi=\theta-\mu d\theta_{k-1}$ and note that (25) can be recast as
$\phi_{k}=\underset{\phi}{\operatorname{argmin}}\;\frac{1}{\lambda}\left\lVert\bar{O}_{k}\phi-\tilde{\epsilon}_{k}\right\rVert^{2}+\left\lVert\phi\right\rVert^{2}.$ (26)
@@
where $\tilde{\epsilon}_{k}=\bar{\epsilon}_{k}-\mu\bar{O}_{k}d\theta_{k-1}$. This formula now has the same form as MinSR with Tikhonov regularization, so we know that the solution is given by
$\phi_{k}=\bar{O}_{k}^{T}(\bar{O}_{k}\bar{O}_{k}^{T}+\lambda I)^{-1}\tilde{\epsilon}_{k}.$ (27)
Finally, we can translate back to $d\theta_{k}$ yielding
$d\theta_{k}=\bar{O}_{k}^{T}(\bar{O}_{k}\bar{O}_{k}^{T}+\lambda I)^{-1}\tilde{\epsilon}_{k}+\mu d\theta_{k-1}.$ (28)
@@
This is almost the formula that we use in practice. The only missing pieces are two extra stabilization procedures that we introduce in Sections 3.1 and 3.2, which are not essential but should be included for best performance. For convenience we present the full procedure for SPRING including these extra details in Algorithm 1. We also include a learning rate schedule $\eta_{k}$ which controls the size of the imaginary time-step used at each iteration.
@@
4:$P\leftarrow\frac{1}{N_{s}}\mathbf{1}\mathbf{1}^{T}$triangleright see Section 3.2
5:$\theta\leftarrow\theta_{0}$
6:$d\theta\leftarrow 0$
7:for k in range(K) do
8:     Sample $R_{1},\ldots,R_{N_{s}}$ from $p(R)=\psi_{\theta}(R)^{2}/\left\langle\psi_{\theta}\middle|\psi_{\theta}\right\rangle$triangleright i.e. using MCMC
9:     $\delta\tau\leftarrow\eta_{k}$
10:     Calculate $\bar{O}$, $\bar{\epsilon}$triangleright via (14)
11:     $\tilde{\epsilon}\leftarrow\bar{\epsilon}-\mu\bar{O}d\theta$
12:     $d\theta\leftarrow\bar{O}^{T}(\bar{O}\bar{O}^{T}+\lambda I+P)^{-1}\tilde{\epsilon}+\mu d\theta$triangleright following (32) and using Cholesky
13:     $d\theta\leftarrow d\theta\cdot\min(1,\sqrt{C}/\left\lVert d\theta\right\rVert)$triangleright norm constraint, see Section 3.1
14:     $\theta\leftarrow\theta+d\theta$
15:end for
16:return $\theta$
It is worth noting that due to several technical details, we cannot directly transfer the convergence results for the classical Kaczmarz method to the VMC setting. First of all, the inclusion of the parameter $\mu$ in SPRING distinguishes it substantially from the Kaczmarz method. We can understand this distinction as arising from the fact that in SPRING, the underlying least-squares system changes a small amount after each iteration. Thus, it makes sense to progressively forget the information from previous iterations, which is what we accomplish by setting $\mu<1$. Additionally, the least-squares problem arising in VMC need not be consistent, and in the inconsistent setting the Kaczmarz method only converges to within a ball of the true solution, with the radius of the ball depending on how inconsistent the system is needell2010randomized . For these reasons, we cannot derive a rigorous convergence guarantee for SPRING with regard to either $d\theta$ or $\theta$. Still, the connection to the Kaczmarz method serves to motivate the algorithm and explain its superior performance relative to MinSR.
@@ 3.1 Norm Constraint
The norm constraint in KFAC is “natural” in the sense that it applies to the norm induced by the Fisher information matrix as opposed to the Euclidean norm. To define this scheme let $g$ be the estimated energy gradient, $\eta$ be the learning rate, $F$ be KFAC’s approximation to the Fisher information matrix, and $d\theta=\eta F^{-1}g$ be the KFAC parameter update. The norm constraint relies upon the fact that $\left\lVert d\theta\right\rVert_{F}^{2}=d\theta^{T}Fd\theta=d\theta^{T}(\eta g)$ and as a result it enforces the constraint
$d\theta^{T}(\eta g)\leq C.$ (29)
If this constraint is not satisfied by the update $d\theta$, then the update is scaled by the unique positive real number $c$ such that $\left\lVert c\cdot d\theta\right\rVert_{F}^{2}=c^{2}d\theta^{T}(\eta g)=C.$
In SPRING the update $d\theta$ and the gradient $g$ are not related by a positive definite matrix F; hence the above scheme is not justified. Instead we implement a simple Euclidean norm constraint given by
@@
Just like in KFAC, if this constraint is violated, we scale the update $d\theta$ by the unique positive real number $c$ such that $\left\lVert c\cdot d\theta\right\rVert^{2}=c^{2}\left\lVert d\theta\right\rVert^{2}=C$. Equivalently, we replace the update $d\theta$ with the scaled update $d\theta\cdot\min(1,\sqrt{C}/\left\lVert d\theta\right\rVert)$.
@@ 3.2 Numerically Stable Inversion
In exact arithmetic, $T=\bar{O}\bar{O}^{T}$ is positive semidefinite, $T+\lambda I$ is positive definite, and $(T+\lambda I)^{-1}\tilde{\epsilon}$ can be computed efficiently using the Cholesky decomposition. However, in single precision we have found that the $T$ matrix can be become indefinite. As a result, the Cholesky decomposition can fail if the regularization is chosen to be too small, and we have even observed this to occur when using our default value of $\lambda=0.001$.
We now note that $T$ is always singular due to the construction of $\bar{O}$. To see this, let $\mathbf{1}$ be the column vector whose entries are all equal to $1$ and let $P=\frac{1}{N_{s}}\mathbf{1}\mathbf{1}^{T}$ represent the orthogonal projector onto the span of $\mathbf{1}$. We can then express $\bar{O}$ as $\bar{O}=\left(I-P\right)O$. As a result, we have that
@@
We alleviate this issue by replacing the matrix $T+\lambda I$ with the further regularized matrix $T+\lambda I+P$. We thus calculate the update for iteration $k$ as
$d\theta_{k}=\bar{O}_{k}^{T}(T_{k}+\lambda I+P)^{-1}\tilde{\epsilon}_{k}+\mu d\theta_{k-1}.$ (32)
This does not affect the value of $d\theta_{k}$ in exact arithmetic. To see this, note that $\mathbf{1}$ is an eigenvector of $(T_{k}+\lambda I)$ with eigenvalue $\lambda$, and is thus also an eigenvector of $(T_{k}+\lambda I+P)$ with eigenvalue $1+\lambda$. As a result, we have
$(T_{k}+\lambda I+P)^{-1}=(T_{k}+\lambda I)^{-1}+\left(\frac{1}{1+\lambda}-\frac{1}{\lambda}\right)P.$ (33)
@@
$\bar{O}_{K}^{T}(T_{k}+\lambda I+P)^{-1}$ $=\bar{O}_{k}^{T}(T_{k}+\lambda I)^{-1}+\left(\frac{1}{1+\lambda}-\frac{1}{\lambda}\right)\bar{O}_{k}^{T}P$
@@
Thus, using (32) reduces the likelihood that numerical errors cause the Cholesky solver to fail without otherwise affecting the computation.
@@ 3.3 MinSR with Momentum
We motivate SPRING primarily via its relation to the Kaczmarz method. An alternative viewpoint is to see SPRING as a modification of a momentum method with an added projection step to remove the component of the momentum that conflicts with the current subproblem $\bar{O}_{k}d\theta=\bar{\epsilon}$. From this perspective, it is natural to wonder how SPRING would compare to MinSR with naïve momentum, which we refer to as MinSR+M. MinSR+M uses the parameter update
$d\theta_{k}=\mu d\theta_{k-1}+(1-\mu)d\phi_{k},$ (34)
where $d\phi_{k}$ represents the MinSR update (without momentum) at iteration $k$, and $\mu$ is the momentum parameter. We reuse the symbol $\mu$ for the momentum parameter since, just like in SPRING, $\mu$ acts as a decay factor on the previous parameter update. We demonstrate in our numerical experiments that SPRING outperforms MinSR+M substantially.

### v2

@@ 3 Methods
Since $R$ is drawn from a continuous and high-dimensional space this equation can be seen as a highly overdetermined linear least-squares problem, with each configuration $R$ corresponding to a single row. Now, let $\bar{O}_{k}$ and $\bar{\epsilon}_{k}$ denote the subsamples from $\bar{O}(R)$ and $\bar{\epsilon}(R)$ that are available at VMC training iteration $k$. Our approach hinges on the assumption that the imaginary time-step is small, meaning the parameter vector $\theta$ changes only slightly at each iteration. Thus, although the sampled subproblems $(\bar{O}_{k-1},\bar{\epsilon}_{k-1})$, $(\bar{O}_{k},\bar{\epsilon}_{k})$ can be completely different, the underlying SR equations are nearly identical. This inspires us to proceed as follows:
@@
    $\underset{d\theta^{\prime}}{\operatorname{argmin}}\;\mathbb{E}_{R\sim p}\left[\left|\bar{O}(R)d\theta^{\prime}-\bar{\epsilon}(R)\right|^{2}\right].$ (26)
@@
    Apply the randomized block Kazcmarz method to these iterates to yield an approximate solution $\phi_{k}$ to the current SR equation.
  3. 3.
    Apply some scaling and clipping procedures, to be described later, to $\phi_{k}$ to calculate the parameter update $d\theta_{k}$.
@@
Following the structure of a single step of the randomized block Kaczmarz method [29], we calculate $\phi_{k}$ by projecting $\phi_{k-1}$ onto the solution space of the underdetermined equation $\bar{O}_{k}\phi=\bar{\epsilon}_{k}$. Recall that the basic form for the Kaczmarz projection is
$\phi_{k}=\phi_{k-1}+\bar{O}_{k}^{+}(\bar{\epsilon}_{k}-\bar{O}_{k}\phi_{k-1}).$ (27)
Equivalently we can say that $\phi_{k}$ satisfies
$\phi_{k}=\operatorname*{arg\,min}_{\phi}\left\lVert\phi-\phi_{k-1}\right\rVert^{2}\text{ s.t. }\bar{O}_{k}\phi=\bar{\epsilon}_{k}.$ (28)
Thus, at a conceptual level, our algorithm differs from MinSR only in how we break the indeterminacy of the sampled subproblem: MinSR chooses the solution of minimal norm, whereas SPRING chooses the solution that is nearest to the previous approximate solution.
In practice $\bar{O}_{k}$ can be singular or very ill-conditioned, so it is beneficial add some form of regularization to this projection. We choose to regularize the projection by incorporating the linear equations using a penalty term rather than a hard constraint. To make this modification, we introduce a regularization parameter $\lambda$ and use the formula
@@
$\phi_{k}=\underset{\phi}{\operatorname{argmin}}\;\frac{1}{\lambda}\left\lVert\bar{O}_{k}\phi-\bar{\epsilon}_{k}\right\rVert^{2}+\left\lVert\phi-\phi_{k-1}\right\rVert^{2}.$ (29)
@@
Even with this regularization, we find that directly using (29) can result in unstable optimization trajectories (see Figure 10(b)). We do not yet have an explanation for this phenomenon. To stabilize the method, we decay the previous gradient by a small amount before projecting it. Formally, this stabilization comes in the form of a new regularization parameter $\mu$ and a modified update formula
@@
$\phi_{k}=\underset{\phi}{\operatorname{argmin}}\;\frac{1}{\lambda}\left\lVert\bar{O}_{k}\phi-\bar{\epsilon}_{k}\right\rVert^{2}+\left\lVert\phi-\mu\phi_{k-1}\right\rVert^{2}.$ (30)
@@
An explicit formula for $\phi_{k}$ can now be derived. We first define $\pi_{k}=\phi_{k}-\mu\phi_{k-1}$, $\pi=\phi-\mu\phi_{k-1}$, and $\bar{\zeta}_{k}=\bar{\epsilon}_{k}-\mu\bar{O}_{k}\phi_{k-1}$ and note that (30) can be recast as
[…]
@@
This formula now has the same form as MinSR with Tikhonov regularization, so we know that the solution is given by
$\pi_{k}=\bar{O}_{k}^{T}(\bar{O}_{k}\bar{O}_{k}^{T}+\lambda I)^{-1}\bar{\zeta}_{k}.$ (32)
Finally, we can translate back to $\phi_{k}$ yielding
$\phi_{k}=\bar{O}_{k}^{T}(\bar{O}_{k}\bar{O}_{k}^{T}+\lambda I)^{-1}\bar{\zeta}_{k}+\mu\phi_{k-1}.$ (33)
To understand this formula, it is helpful to explicitly write out the three terms:
$\phi_{k}=\bar{O}_{k}^{T}(\bar{O}_{k}\bar{O}_{k}^{T}+\lambda I)^{-1}\bar{\epsilon}_{k}+\mu\phi_{k-1}-\mu\bar{O}_{k}^{T}(\bar{O}_{k}\bar{O}_{k}^{T}+\lambda I)^{-1}\bar{O}_{k}\phi_{k-1}.$ (34)
The first term alone corresponds to MinSR, while the first two terms together correspond to a scaled version of MinSR with momentum (to be presented in Section 3.4). It is the addition of the third term that concretely distinguishes SPRING from these other methods.
@@
Equation (33) is almost the formula that we use in practice. The only missing piece is an extra stabilization procedure that we introduce in Section 3.2, which is not essential but should be included for best performance. Finally, to calculate $d\theta_{k}$, we also include a learning rate schedule $\eta_{k}$ and a norm constraint $C$ which ensures that the parameters are not changed too much at each iteration (see Section 3.1). To ensure that the Kaczmarz algorithm is implemented consistently, regardless of the learning rate schedule, we always calculate $\bar{\epsilon}_{k}$ using $\delta\tau=1$ and then apply the learning rate $\eta_{k}$ only when we calculate $d\theta_{k}$. For convenience we present the full procedure for SPRING including these extra details in Algorithm 1.
@@
4:$\theta\leftarrow\theta_{0}$
5:${\phi}\leftarrow 0$
6:for k in 1:K do
7:     Sample $R_{1},\ldots,R_{N_{s}}$ from $p(R)=\psi_{\theta}(R)^{2}/\left\langle\psi_{\theta}\middle|\psi_{\theta}\right\rangle$triangleright i.e. using MCMC
8:     Calculate $\bar{O}$, $\bar{\epsilon}$ using $\delta\tau=1$triangleright via (13), (14)
9:     $\bar{\zeta}\leftarrow\bar{\epsilon}-\mu\bar{O}\phi$
10:     $\phi\leftarrow\bar{O}^{T}(\bar{O}\bar{O}^{T}+\lambda I+\frac{1}{N_{s}}\mathbf{1}\mathbf{1}^{T})^{-1}\bar{\zeta}+\mu\phi$triangleright following (39) and using Cholesky
11:     $d\theta\leftarrow\phi\cdot\min(\eta_{k},\sqrt{C}/\left\lVert\phi\right\rVert)$triangleright See Section 3.1
12:     $\theta\leftarrow\theta+d\theta$
13:end for
14:return $\theta$
It is worth noting that due to several technical details, we cannot directly transfer the convergence results for the classical Kaczmarz method to the VMC setting. First of all, the inclusion of the parameter $\mu$ in SPRING distinguishes it substantially from the Kaczmarz method. We can understand this distinction as arising from the fact that in SPRING, the underlying least-squares system changes a small amount after each iteration. Thus, it makes sense to progressively forget the information from previous iterations, which is what we accomplish by setting $\mu<1$. Additionally, the least-squares problem arising in VMC need not be consistent, and in the inconsistent setting the Kaczmarz method only converges to within a ball of the true solution, with the radius of the ball depending on how inconsistent the system is [38]. For these reasons, we cannot derive a rigorous convergence guarantee for SPRING with regard to either $d\theta$ or $\theta$. Still, the connection to the Kaczmarz method serves to motivate the algorithm and explain its superior performance relative to MinSR.
@@ 3.1 Norm Constraint
The norm constraint in KFAC is “natural” in the sense that it applies to the norm induced by the Fisher information matrix as opposed to the Euclidean norm. To define this scheme let $g$ be the estimated energy gradient, $\eta$ be the learning rate, $S$ be KFAC’s approximation to the Fisher information matrix, and $\phi=S^{-1}g$ be the unscaled and unconstrained KFAC parameter update. The final update will be $d\theta=c\phi$ for some scalar $c$. The norm constraint relies upon the fact that $\left\lVert d\theta\right\rVert_{S}^{2}=c^{2}\phi^{T}S\phi=c^{2}\phi^{T}g$ and as a result it enforces the constraint
$c^{2}\phi^{T}g\leq C.$ (35)
Concretely this is achieved by choosing
$d\theta=\phi\cdot\min(\eta,\sqrt{C}/\sqrt{\phi^{T}g}).$ (36)
In SPRING the update $d\theta$ and the gradient $g$ are not related by a positive definite matrix S; hence the above scheme is not justified. Instead we implement a simple Euclidean norm constraint given by
@@
Just like in KFAC, if this constraint is violated before the norm constraint is applied, we scale the update down to ensure that it is satisfied. Concretely, we calculate the update $d\theta_{k}$ from the Kaczmarz iterate $\phi_{k}$ and the learning rate $\eta_{k}$ using the formula
$d\theta_{k}=\phi_{k}\cdot\min(\eta_{k},\sqrt{C}/\left\lVert\phi_{k}\right\rVert).$
@@ 3.2 Numerically Stable Inversion
In exact arithmetic, $T=\bar{O}\bar{O}^{T}$ is positive semidefinite, $T+\lambda I$ is positive definite, and $(T+\lambda I)^{-1}\bar{\zeta}$ can be computed efficiently using the Cholesky decomposition. However, in single precision we have found that the $T$ matrix can be become indefinite. As a result, the Cholesky decomposition can fail if the regularization is chosen to be too small, and we have even observed this to occur when using our default value of $\lambda=0.001$.
We now note that $T$ is always singular due to the construction of $\bar{O}$. Recall that $\mathbf{1}$ represents the column vector whose entries are all equal to $1$ and $P=\frac{1}{N_{s}}\mathbf{1}\mathbf{1}^{T}$ represents the orthogonal projector onto the span of $\mathbf{1}$. Since $\bar{O}=\left(I-P\right)O$, we have that
@@
We alleviate this issue by replacing the matrix $T+\lambda I$ with the further regularized matrix $T+\lambda I+\omega P$ for some positive real $\omega$. We thus calculate the Kaczmarz iterate for iteration $k$ as
$\phi_{k}=\bar{O}_{k}^{T}(T_{k}+\lambda I+\omega P)^{-1}\bar{\zeta}_{k}+\mu\phi_{k-1}.$ (39)
This does not affect the value of $\phi_{k}$ in exact arithmetic. To see this, note that $\mathbf{1}$ is an eigenvector of $(T_{k}+\lambda I)$ with eigenvalue $\lambda$, and is thus also an eigenvector of $(T_{k}+\lambda I+\omega P)$ with eigenvalue $\omega+\lambda$. As a result, we have
$(T_{k}+\lambda I+\omega P)^{-1}=(T_{k}+\lambda I)^{-1}+\left(\frac{1}{\omega+\lambda}-\frac{1}{\lambda}\right)P.$ (40)
@@
$\bar{O}_{K}^{T}(T_{k}+\lambda I+\omega P)^{-1}$ $=\bar{O}_{k}^{T}(T_{k}+\lambda I)^{-1}+\left(\frac{1}{\omega+\lambda}-\frac{1}{\lambda}\right)\bar{O}_{k}^{T}P$
@@
Thus, using (39) reduces the likelihood that numerical errors cause the Cholesky solver to fail without otherwise affecting the computation. In our experiments we have used the value $\omega=1$ in all cases. However, it is theoretically possible for the eigenvalues of $T$ to all be smaller than $1$, in which case setting $\omega=1$ would increase the condition number of the regularized $T$ matrix. In such a case $\omega$ could instead be chosen for example to be the mean of the eigenvalues of the original $T$ matrix.
3.3 MinSR implementation details
We now briefly discuss the version of MinSR that we use for our numerical experiments, which is very similar but not identical to the basic version presented in Section 2.3. Functionally, the only differences are the inclusion of a norm constraint and the extra regularization of the matrix $\bar{O}_{k}\bar{O}_{k}^{T}$ as described in Section 3.2. Additionally, for consistency with SPRING we set $\delta\tau=1$ when calculating $\bar{\epsilon}_{k}$ and we instead control the step size with a learning rate parameter $\eta_{k}$. With these modifications the MinSR algorithm is described by the following formulas:
$\phi_{k}=\bar{O}_{k}^{T}(\bar{O}_{k}\bar{O}_{k}^{T}+\lambda I+P)^{-1}\bar{\epsilon}_{k},$ (41)
$d\theta_{k}=\phi_{k}\cdot\min(\eta_{k},\sqrt{C}/\left\lVert\phi_{k}\right\rVert).$ (42)
@@ 3.4 MinSR with Momentum
We motivate SPRING primarily via its relation to the Kaczmarz method. An alternative viewpoint is to see SPRING as a modification of a momentum method with an added projection step to remove the component of the momentum that conflicts with the current subproblem $\bar{O}_{k}d\theta=\bar{\epsilon}$. While we do not endorse this perspective, it is natural to wonder how SPRING would compare to MinSR with naïve momentum, which we refer to as MinSR+M.
Concretely, MinSR+M iteratively updates a momentum direction $\phi_{k}$ by taking a convex combination of the previous value $\phi_{k-1}$ and the current MinSR solution. Just like in SPRING, we assume an imaginary time step of $\delta\tau=1$ when we calculate $\bar{\epsilon}_{k}$ and instead control the step size using a learning rate parameter $\eta_{k}$ and a norm constraint. The method is thus described by the formulas:
$\phi_{k}=(1-\mu)\bar{O}_{k}^{T}(\bar{O}_{k}\bar{O}_{k}^{T}+\lambda I+P)^{-1}\bar{\epsilon}_{k}+\mu\phi_{k-1},$ (43)
$d\theta_{k}=\phi_{k}\cdot\min(\eta_{k},\sqrt{C}/\left\lVert\phi_{k}\right\rVert).$ (44)
As in SPRING, the $\mu$ parameter acts as a decay factor on the previous parameter update. However, MinSR+M differs significantly from SPRING in that no regularized projector is applied to the previous update. It also differs in that the MinSR solution is scaled by a factor of $(1-\mu)$ to achieve a convex combination, which means $\mu$ must be chosen to be strictly less than one. See Section 4.3 and Figure 10(a) for further discussion of the choice of $\mu$ for MinSR+M. We demonstrate in our numerical experiments that SPRING outperforms MinSR+M substantially. This supports the viewpoint that SPRING is not simply a momentum method, but is rather a way of improving the quality of the parameter updates as approximate solutions to the SR equation.